A Boeing 747 “Jumbo Jet” has a length of 59.7 m. The runway on which the plane lands intersects another runway. The width of the intersection is 25.0 m. The plane accelerates through the intersection at a rate of -5.5 m/s2 and clears it with a final speed of 47.9 m/s. How much time is needed for the plane to clear the intersection? Use 3 significant figures in your final answer.
Answer #1

here,

length of Jumbo Jet, l = 59.7 m

the width of intersection , w = 25 m

acceleration , a = - 5.5 m/s^2

final speed . v = 47.9 m/s

let the initial speed be u and time taken be t

using third equation of motion

v^2 - u^2 =2 * a * ( l + w)

47.9^2 - u^2 = 2 * (-5.5) * ( 59.7 + 25)

solving for u

u = 58.3 m/s

using first equation of motion

v = u + a * t

47.9 = 58.3 - 5.5 * t

solving for t

t = 1.90 s

the time taken is 1.90 s